Given the Concentration (M) being 5.39x10^-5 M, what is the concentration in mg/mL?

0.25 = 2187x + 0.1322 (0.25 – 0.1322) = 2187x .1178 = = X 2187 x = 5.386374028x10-5 ~5.39x10-5M
Answer #1

5.39*10^-5 M means 1 liter (1000 ml) solution contains 5.39*10^-5 moles of solute.

Let X g/mol is the molar mass of solute. Then,

Mass of solute = moles * molar mass = (5.39*10^-5)*X grams

= 5.39*10^-5 * X * 1000 miligrams = (5.39*10^-2)X mg

Now,

Concentration in mg/ml = (5.39*10^-2)*X / 1000 = (5.39*10^-5)*X mg/ml .....Answer

Let me know if any doubts.